# Test time training enhances in-context learning of nonlinear functions

### v1

@@ 2.4 Training algorithm
  • TTT stage I: We apply a single gradient descent step to ${\bm{u}}^{(0)}$ with $L_{2}$-regularization, using $({\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}})$ as contexts and considering the output from the original model as a teacher signal. This prevents catastrophic forgetting, a phenomenon where the parameter change caused by TTT deprives the model of the fundamental ability to adapt to the original task. The necessity of this stage is further discussed in Subsection 3.1.2. The goal of this stage is to find a good initial value of ${\bm{u}}^{(1)}$ that satisfies $\left\langle\beta,{\bm{u}}^{(1)}\right\rangle\geq 1/\mathrm{polylog}(d)$.
  • TTT stage II: We continue to optimize $\bm{u}$} by applying multi-step online SGD scheme that originates from Lee et al. (2024), using $({\bm{X}}_{N_{2}},{\bm{y}}_{N_{2}})$ as contexts. This stage aims to align $\bm{u}$} more closely to the target $\beta$. As we will see, increasing the number of the SGD steps leads to the convergence of $\bm{u}$} to $\beta$.
@@
1
@@
2
3Initialize $\Gamma(0)\sim{\bm{I}}_{d}/\sqrt{d},~{\bm{v}}(0)\sim\mathrm{Unif}(\{\pm 1\}^{m}),~{\bm{b}}(0)=\bm{0}_{m},~{\bm{a}}(0)=\alpha_{pt}\bm{1}_{m}$.
4Pretraining: One-step Gradient descent on Attention Matrix
5 $\Gamma^{*}\leftarrow\Gamma{(0)}-\eta_{pt}\frac{1}{2T_{pt}}\sum_{t=1}^{T_{pt}}\nabla_{\Gamma}((f_{\mathrm{IC}}({\bm{x}}^{t})-y^{t})^{2}+\lambda_{pt}\|\Gamma\|_{F}^{2})$.
6
7Test-Time Training
8 ${\bm{a}}(0)=\alpha_{1}\bm{1}_{m}$.
9 for $i=1$ to $N_{1}$ do
10    ${\bm{x}}_{i}\leftarrow\sqrt{r}\Gamma^{*}{\bm{x}}_{i}$.
11 ${\bm{u}}^{(0)}\sim\mathcal{N}(0,{\bm{I}}_{d})$, ${\bm{u}}^{(0)}\leftarrow\sqrt{r}\Gamma_{*}{\bm{u}}^{(0)}$, ${\bm{u}}^{(0)}\leftarrow\frac{{\bm{u}}^{(0)}}{\sqrt{r}\|{\bm{u}}^{(0)}\|}$.
12
13 Stage I: Initialization of u with signals from Original Model
14      Draw ${\bm{w}}_{1},\cdots,{\bm{w}}_{N_{new}}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d})$.
15    $b=\frac{1}{N_{new}}\sum_{i=1}^{N_{new}}g(\Gamma^{*},{\bm{w}}_{i})$,
16    ${\bm{u}}^{(1)}={\bm{u}}^{(0)}-\eta_{1}\{\sqrt{r}\Gamma^{*}\nabla_{{\bm{u}}}\frac{1}{2N_{new}}\sum_{i=1}^{N_{new}}(f_{\mathrm{IC}}({\bm{w}}_{i})-(g(\Gamma_{*},{\bm{w}}_{i})-b))^{2}+\frac{\lambda_{1}}{2}\nabla_{{\bm{u}}}\|{\bm{u}}\|^{2}\}$,
17    ${\bm{u}}^{(1)}\leftarrow\frac{{\bm{u}}^{(1)}}{\|{\bm{u}}^{(1)}\|}$.
18
19 ${\bm{a}}(0)=\alpha_{2}\bm{1}_{m}$.
20 for $i=N_{1}+1$ to $N_{1}+N_{2}$ do
21    ${\bm{x}}_{i}\leftarrow\sqrt{r}\Gamma^{*}{\bm{x}}_{i}$.
22 Stage II: Strong Recovery
23    for $t=1$ to $t=N_{3}$ do
24       ${\bm{u}}^{(t+1)}={\bm{u}}^{(t)}-\eta_{2}\sqrt{r}\Gamma^{*}\nabla_{u}(\frac{1}{2}(f_{\mathrm{IC}}({\bm{x}}_{N_{1}+N_{2}+t})-y_{N_{1}+N_{2}+t})^{2})$. ${\bm{u}}^{(t+1)}\leftarrow\frac{{\bm{u}}^{(t+1)}}{\|{\bm{u}}^{(t+1)}\|}$.
25
26
27 Initialize $b_{j}^{*}\sim\mathrm{Unif}([-\log^{2}{d},\log^{2}{d}])$ , ${\bm{v}}^{*}={\bm{v}}(0)$.
28 Stage III: Training of MLP Layer
29    ${\bm{a}}^{*}\leftarrow\arg\min_{{\bm{a}}}\frac{1}{2N_{4}}\sum_{t=N_{1}+N_{2}+N_{3}+1}^{N_{1}+N_{2}+N_{3}+N_{4}}(f_{\mathrm{TF}}({\bm{x}}_{t},{\bm{u}}^{(N_{3}+1)},{\bm{v}}^{*},{\bm{a}},{\bm{b}}^{*})-y_{t})^{2}+\frac{\lambda_{2}}{2}\|{\bm{a}}\|^{2}$.
30
31
@@ 28 Stage III: Training of MLP Layer
Algorithm 1 Pretraining and test-time training of transformer
@@ 3 Main result
Theorem 1 (Formal).
We denote the link function drawn in inference-time as $\sigma_{*}^{\mathrm{test}}$. Suppose that $T_{pt},N_{pt}=\tilde{\Omega}(r^{2}d^{Q+2})$, $N_{1}=\tilde{\Omega}(r^{\mathrm{ge}(\sigma_{*}^{\mathrm{test}})+2})$, $N_{new}=\tilde{\Omega}(r^{\mathrm{ge}(\sigma_{*}^{\mathrm{test}})+2})$, $N_{2}=\tilde{\Theta}(r^{2})$. Moreover, we assume that $m,N_{pt},T_{pt},N_{1},N_{2},N_{3},N_{4}=O(\mathrm{poly}(d))$ and there exists $\alpha_{r}>0$ that $r=\Omega(d^{\alpha_{r}})$. When we fix $d$ large enough, then there exists $\lambda_{pt},\lambda_{1},\eta_{pt},\eta_{1},\eta_{2}$ such that the prediction risk is low with probability at least 0.99 over the training data and random initialization. Concretely, for the model trained via Algorithm 1, we have that
$|\mathcal{R}_{f_{\mathrm{TF}}}({\bm{u}},{\bm{v}}^{*},{\bm{a}}^{*},{\bm{b}}^{*})-\tau|=\tilde{O}(N_{4}^{-1/2})+\tilde{O}(m^{-1/2})+\tilde{O}\Bigg(\sqrt{\frac{r\sqrt{r}}{N_{3}}}\Bigg),$
@@
(ii)The test-time sample complexity
Theorem 1 implies that $N_{test}=\tilde{\Omega}(r^{2+\mathrm{ge}(\sigma_{*}^{\mathrm{test}})})$ to ensure low predictive loss $o(1)$. While this is slightly looser than the $\tilde{\Omega}(r^{3\mathrm{ge}(\sigma_{*})/2})$ presented in Nishikawa et al. (2025), our result still has two key benefits. First, our sample complexity does not depend on the entire dimension $d$, which means this model also adapts to the low-dimensionality of $\beta$. Second, our sample complexity is independent of either $\mathrm{deg}(\sigma_{*}^{\mathrm{test}})$ or $\mathrm{ie}(\sigma_{*}^{\mathrm{test}})$, which breaks the CSQ upper bound. In addition, for polynomial $\mathrm{ge}(\sigma_{*})\leq 2$ holds, so the required number of samples remains modest. In conclusion, our sample complexity is highly competitive with that of prior work.
(iii) The adaptability to link function
Our framework has the flexibility of using a task-specific link function $\sigma_{*}^{t}$. In the previous work by Nishikawa et al. (2025), the link function is fixed throughout all the tasks because the training of the relevant layer is done only in pretraining. In contrast, the use of test-time training allows the model to adapt to the characteristics of each task. Therefore, our algorithm is effective even when the underlying link function varies from one task to another.
(iv) The convergence rate of predictive loss with respect to n
The analysis by Nishikawa et al. (2025) guarantees that the ICL risk is $o_{d}(1)$, but it provides no guarantee that the risk diminishes as $n\rightarrow\infty$. In fact, this is an inherent consequence of adopting softmax attention, as we discussed in Remark 4. In contrast, our result overcomes that limitation. When $d$ is a sufficiently large constant, our theory ensures that the prediction risk can be made arbitrarily close to the inevitable noise $\tau$ by increasing the context size $N$ and the network width $m$. This is because the increase in the test-time context length $N_{3}$ allows the vector $\bm{u}$} to converge to the ground truth $\beta$.
@@ 3.1 Proof sketch
The outline of the proof of our main theorem is as follows: using the output of the original model, we can achieve weak recovery i.e. nontrivial overlap between $\beta$ and $\bm{u}$}. Once weak recovery is completed, more training using the signal $y$ leads to strong recovery i.e. $\left\langle\beta,{\bm{u}}\right\rangle\geq 1-\varepsilon$. Finally, the training of MLP layer allows the model to fit to the nonlinearity of the link function.
@@ 3.1.1 Exploiting the pretrained attention matrix
Lemma 5 (Informal, Correspond to Proposition 22 in Nishikawa et al. (2025)).
@@
This means that $\Gamma^{*}$ can project vectors into the $r$-dimensional subspace $\mathrm{Supp}(\beta)$. Therefore, by multiplying $\sqrt{r}\Gamma^{*}$ to ${\bm{x}}_{i}$ and the gradient (the coefficient $\sqrt{r}$ is just for adjusting the scale), we can make the problem virtually $r$-dimensional, even though the entire dimension is $d$. This leads to the sample complexity only scaling up with $r$, not the whole dimension $d$.
@@ 3.1.2 Weak recovery
First, we initialize $\bm{u}$} by one step gradient descent using the output from the original model $f_{\mathrm{IC}}({\bm{w}}_{i})$. We do not use the signal $y$ in this process. Intuitively, this self-distillation can be seen as a prevention of catastrophic forgetting. Taking the information from the original model makes the LoRA model similar to the original model, thereby preserving the desired features of the original model.
@@
$\frac{1}{2}\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{(f_{\mathrm{IC}}({\bm{x}})-y)^{2}}\approx\tilde{\Theta}(\alpha m)\left\langle\beta,{\bm{u}}\right\rangle\{(\sqrt{r})^{-(\mathrm{ie}(\sigma_{*})-1)}+\left\langle\beta,{\bm{u}}\right\rangle^{2\mathrm{ie}(\sigma_{*})-2}\}\beta.$
@@ 3.1.3 Strong recovery
Weak recovery is insufficient for reliably predicting $\left\langle\beta,{\bm{x}}\right\rangle$ using only $\bm{u}$}. Therefore, we further optimize $\bm{u}$} until we achieve strong recovery, defined as $\left\langle\beta,{\bm{u}}\right\rangle\geq 1-\varepsilon$. After achieving $\left\langle\beta,{\bm{u}}^{(1)}\right\rangle>1/\mathrm{polylog}(d)$, the signal strength is $O(1/\mathrm{polylog}(d))$, independent of $\mathrm{ge}(\sigma_{*}^{\mathrm{test}})$. Consequently, the data length required to achieve strong recovery does not depend on $\mathrm{ge}(\sigma_{*}^{\mathrm{test}})$. Moreover, we achieved a smaller sample size for achieving strong recovery compared to the prior work by Lee et al. (2024). While their work suggested a linear increase in $\left\langle\beta,{\bm{u}}^{(n)}\right\rangle$, which resulted in the required data length $N=\Theta_{\varepsilon}(\varepsilon^{-2})$, we find that beyond a certain point, $1-\left\langle\beta,{\bm{u}}^{(n)}\right\rangle$ converges to 0 geometrically. This accelerated geometric convergence yields an improved sample complexity of $N=\Theta_{\varepsilon}(\frac{1}{\varepsilon}\log{\frac{1}{\varepsilon}})$. See Appendix E for further details.

### v2

@@ 2.4 Training algorithm
  • TTT stage I: We apply a single gradient descent step to ${\bm{u}}^{(0)}$ with $L_{2}$-regularization. This stage prevents catastrophic forgetting, a phenomenon where the parameter change caused by TTT deprives the model of the fundamental ability to adapt to the original task. The goal of this stage is to find a good initial value of ${\bm{u}}^{(1)}$ that satisfies $\left\langle\beta,{\bm{u}}^{(1)}\right\rangle\geq 1/\mathrm{polylog}(d)$. For this stage, we use the output from the attention layer of the original model as a teacher signal, which is defined as $g(\Gamma_{*},{\bm{X}}_{N1},{\bm{y}}_{N_{1}},{\bm{x}})=\frac{N_{1}^{-1}\sum_{i=1}^{N_{1}}y_{i}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma_{*}{\bm{x}}}}{N_{1}^{-1}\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma_{*}{\bm{x}}}}$. In this stage, the query $\bm{x}$} do not require ground-truth label $y$. Therefore, we use $N_{new}$ newly generated vectors $w_{1},\dots,w_{N_{new}}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d})$ for the query. The necessity of this stage is further discussed in subsection 3.1.2.
  • TTT stage II: We continue to optimize $\bm{u}$} by applying multi-step online SGD scheme that originates from Lee et al. (2024), using $({\bm{X}}_{N_{2}},{\bm{y}}_{N_{2}})$ as contexts. In this stage, we use the ground truth $({\bm{X}}_{N_{3}},{\bm{y}}_{N_{3}})$ as the teacher signal. This stage aims to align $\bm{u}$} more closely to the target $\beta$. As we will see, increasing the number of the SGD steps leads to the convergence of $\bm{u}$} to $\beta$.
@@
Algorithm 1 Pretraining and test-time training of transformer
@@
$\Gamma(0)\sim{\bm{I}}_{d}/\sqrt{d},~{\bm{v}}(0)\sim\mathrm{Unif}(\{\pm 1\}^{m}),~{\bm{b}}(0)=\bm{0}_{m},~{\bm{a}}(0)=\alpha_{pt}\bm{1}_{m}$.
{Pretraining}
$\Gamma^{*}\leftarrow\Gamma{(0)}-\eta_{pt}\frac{1}{2T_{pt}}\sum_{t=1}^{T_{pt}}\nabla_{\Gamma}((f_{\mathrm{IC}}(\Gamma,{\bm{X}}_{N_{pt}},{\bm{y}}_{N_{pt}},{\bm{x}}^{t})-y^{t})^{2}+\lambda_{pt}\|\Gamma\|_{F}^{2})$.
 {Test-time Training}
${\bm{a}}(0)=\alpha_{1}\bm{1}_{m}$.
for $i=1$ to $N_{1}$ do
  ${\bm{x}}_{i}\leftarrow\sqrt{r}\Gamma^{*}{\bm{x}}_{i}$.
end for
${\bm{u}}^{(0)}\sim\mathcal{N}(0,{\bm{I}}_{d})$, ${\bm{u}}^{(0)}\leftarrow\sqrt{r}\Gamma_{*}{\bm{u}}^{(0)}$, ${\bm{u}}^{(0)}\leftarrow\frac{{\bm{u}}^{(0)}}{\sqrt{r}\|{\bm{u}}^{(0)}\|}$.
{Stage I: Weak Recovery}
 Draw ${\bm{w}}_{1},\cdots,{\bm{w}}_{N_{new}}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d})$.
$b=\frac{1}{N_{new}}\sum_{i=1}^{N_{new}}g(\Gamma^{*},{\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},{\bm{w}}_{i})$,
${\bm{u}}^{(1)}={\bm{u}}^{(0)}-\eta_{1}\{\sqrt{r}\Gamma^{*}\nabla_{{\bm{u}}}\frac{1}{2N_{new}}\sum_{i=1}^{N_{new}}$
$\quad(f_{\mathrm{IC}}(\Gamma_{\bm{u}},{\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},{\bm{w}}_{i})$
$\quad-(g(\Gamma_{*},{\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},{\bm{w}}_{i})-b))^{2}+\frac{\lambda_{1}}{2}\nabla_{{\bm{u}}}\|{\bm{u}}\|^{2}\}$,
${\bm{u}}^{(1)}\leftarrow\frac{{\bm{u}}^{(1)}}{\|{\bm{u}}^{(1)}\|}$.
${\bm{a}}(0)=\alpha_{2}\bm{1}_{m}$.
for $i=N_{1}+1$ to $N_{1}+N_{2}$ do
  ${\bm{x}}_{i}\leftarrow\sqrt{r}\Gamma^{*}{\bm{x}}_{i}$.
end for
{Stage II: Strong Recovery}
for $t=1$ to $N_{3}$ do
  ${\bm{u}}^{(t+1)}={\bm{u}}^{(t)}-\eta_{2}\sqrt{r}\Gamma^{*}\nabla_{u}$
  $\quad(\frac{1}{2}(f_{\mathrm{IC}}(\Gamma_{u},{\bm{X}}_{N_{2}},{\bm{y}}_{N_{2}},{\bm{x}}_{N_{1}+N_{2}+t})-y_{N_{1}+N_{2}+t})^{2})$.
  ${\bm{u}}^{(t+1)}\leftarrow\frac{{\bm{u}}^{(t+1)}}{\|{\bm{u}}^{(t+1)}\|}$.
end for
$b_{j}^{*}\sim\mathrm{Unif}([-\log^{2}{d},\log^{2}{d}])$ , ${\bm{v}}^{*}={\bm{v}}(0)$.
{Stage III: Training of MLP Layer}
${\bm{a}}^{*}\leftarrow\arg\min_{{\bm{a}}}\frac{1}{2N_{4}}\sum_{t=N_{1}+N_{2}+N_{3}+1}^{N_{1}+N_{2}+N_{3}+N_{4}}$
$\quad(f_{\mathrm{TF}}({\bm{x}}_{t},{\bm{u}}^{(N_{3}+1)},{\bm{v}}^{*},{\bm{a}},{\bm{b}}^{*})-y_{t})^{2}+\frac{\lambda_{2}}{2}\|{\bm{a}}\|^{2}$.
@@
Remark 2.4.
While Algorithm 1 involves multiple stages, this structure reflects distinct practical objectives rather than arbitrary complexity. In real-world implementations, pretraining and test-time training naturally operate as separate phases. Furthermore, Stage I utilizes a self-supervised objective for initialization, which fundamentally differs from the supervised objectives in the later stages; thus, it must be treated as a distinct step even in practice. The only structural deviation from a typical implementation is the separation of Stage II (Attention learning) and Stage III (MLP learning). This layer-wise decoupling is primarily introduced for theoretical analysis: it is a standard technique for obtaining rigorous guarantees and has been widely adopted in prior works (Damian et al., 2022; Abbe et al., 2023; Lee et al., 2024; Oko et al., 2024).
@@ 3 Main result
Theorem 1.1 (Formal).
We denote the link function drawn in inference-time as $\sigma_{*}^{\mathrm{test}}$. Suppose that $T_{pt},N_{pt}=\tilde{\Omega}(r^{2}d^{Q+2})$, $N_{1}=\tilde{\Omega}(r^{\mathrm{ge}(\sigma_{*}^{\mathrm{test}})+2})$, $N_{new}=\tilde{\Omega}(r^{\mathrm{ge}(\sigma_{*}^{\mathrm{test}})+2})$, $N_{2}=\tilde{\Theta}(r^{2})$. Moreover, we assume that $m,N_{pt},T_{pt},N_{1},N_{2},N_{3},N_{4}=O(\mathrm{poly}(d))$ and there exists $\alpha_{r}>0$ that satisfies $r=\Omega(d^{\alpha_{r}})$. When we fix $d$ large enough, then there exists $\lambda_{pt},\lambda_{1},\eta_{pt},\eta_{1},\eta_{2}$ such that the prediction risk is low with probability at least 0.99 over the training data and random initialization. Concretely, for the model trained via Algorithm 1, we have that
$|\mathcal{R}_{f_{\mathrm{TF}}}({\bm{u}},{\bm{v}}^{*},{\bm{a}}^{*},{\bm{b}}^{*})-\tau|$
$=$ $\tilde{O}(N_{4}^{-1/2})+\tilde{O}(m^{-1/2})+\tilde{O}\Bigg(\sqrt{\frac{r\sqrt{r}}{N_{3}}}\Bigg),$
@@
(ii) The adaptability to link function
Our framework has the flexibility of using a task-specific link function $\sigma_{*}^{t}$. In the previous work by Nishikawa et al. (2025), the link function is fixed throughout all the tasks (only $\beta$ varies across different tasks) because the training of the relevant layer is done only in pretraining. In contrast, the use of test-time training allows the model to adapt to the characteristics of each task. Therefore, our algorithm is effective even when the underlying link function varies from one task to another.
(iii) The convergence rate of predictive loss with respect to n
The analysis by Nishikawa et al. (2025) guarantees that the ICL risk is $o_{d}(1)$, but it provides no guarantee that the risk diminishes as $n\rightarrow\infty$. In fact, this is an inherent consequence of adopting softmax attention, as we discussed in Remark 2.3. In contrast, our result overcomes that limitation. When $d$ is a sufficiently large constant, our theory ensures that the prediction risk can be made arbitrarily close to the inevitable noise $\tau$ by increasing the context size $N_{test}$ and the network width $m$. This is because the increase in the test-time context length $N_{3}$ allows the vector $\bm{u}$} to converge to the ground truth $\beta$.
In addition, it is worth noting that we achieve efficient sample complexity. Theorem 1.1 implies that $N_{test}=\tilde{\Omega}(r^{2+\mathrm{ge}(\sigma_{*}^{\mathrm{test}})})$ to ensure low predictive loss $\tilde{O}(\sqrt{\frac{r\sqrt{r}}{N_{test}}})$, yielding two key benefits. First, our sample complexity does not depend on the entire dimension $d$, which means this model adapts to the low-dimensionality of $\beta$. Second, our sample complexity is independent of either $\mathrm{deg}(\sigma_{*}^{\mathrm{test}})$ or $\mathrm{ie}(\sigma_{*}^{\mathrm{test}})$, which breaks the CSQ upper bound. Moreover, for polynomials where $\mathrm{ge}(\sigma_{*})\leq 2$ holds, the required number of samples is small. In conclusion, our sample complexity is nearly optimal. It is instructive to compare this with Nishikawa et al. (2025). Although they also achieve sample complexity with these two features ($N_{test}=\tilde{\Omega}(r^{3\mathrm{ge}(\sigma_{*})/2})$), their result guarantees convergence with respect to $d$, not $N_{test}$. This implies that their result is valid only in the asymptotic limit where $d\to\infty$. Moreover, they do not provide the convergence rate with respect to $d$. In contrast, our result holds for any sufficiently large fixed $d$.
These advantages are further supported by numerical experiments in section  4.
@@ 3.1 Proof sketch
The outline of the proof of our main theorem proceeds in three stages. First, we achieve weak recovery—i.e., a nontrivial overlap between $\beta$ and $\bm{u}$}—by leveraging the output of the pretrained attention layer as a teacher signal. The attention mechanism reduces the information exponent of the link function $\mathrm{ie}(\sigma_{*})$ to its general exponent $\mathrm{ge}(\sigma_{*})$. We exploit this property to significantly accelerate the initial alignment of $\bm{u}$} compared to vanilla SGD. Once weak recovery is established, we show that subsequent training on the ground-truth signal $y$ drives the model toward strong recovery ($\left\langle\beta,{\bm{u}}\right\rangle\geq 1-\varepsilon$). Finally, the optimization of the MLP layer allows the model to adapt to the task-specific nonlinearity of the target link function.
@@ 3.1.1 Exploiting the pretrained attention matrix
Lemma 3.1 (Informal, Correspond to Proposition 22 in Nishikawa et al. (2025)).
@@
This means that $\Gamma^{*}$ can project vectors into the $r$-dimensional subspace $\mathrm{Supp}(\beta)$. Therefore, by multiplying $\sqrt{r}\Gamma^{*}$ to the context ${\bm{X}}_{N_{1}},{\bm{X}}_{N_{2}}$ and the gradient (the coefficient $\sqrt{r}$ is just for adjusting the scale), we can make the problem virtually $r$-dimensional, even though the entire dimension is $d$. This leads to the sample complexity only scaling up with $r$, not the whole dimension $d$.
@@ 3.1.2 Weak recovery
First, we initialize $\bm{u}$} by one step gradient descent using the output from the original model $g(\Gamma_{*},{\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},w_{i})$. We do not use the signal $y$ in this process. Intuitively, this self-distillation can be seen as a prevention of catastrophic forgetting. Taking the information from the original model makes the LoRA model similar to the original model, thereby preserving the desired features of the original model.
@@
$\frac{1}{2}\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{(f_{\mathrm{IC}}(\Gamma,{\bm{X}}_{N},{\bm{y}}_{N},{\bm{x}})-y)^{2}}$
$\approx\tilde{\Theta}(\alpha m)\left\langle\beta,{\bm{u}}\right\rangle\{(\sqrt{r})^{-(\mathrm{ie}(\sigma_{*})-1)}+\left\langle\beta,{\bm{u}}\right\rangle^{2\mathrm{ie}(\sigma_{*})-2}\}\beta.$
@@ 3.1.3 Strong recovery
Weak recovery is insufficient for reliably predicting $\left\langle\beta,{\bm{x}}\right\rangle$ using only $\bm{u}$}. Therefore, we further optimize $\bm{u}$} until we achieve strong recovery, defined as $\left\langle\beta,{\bm{u}}\right\rangle\geq 1-\varepsilon$ for a small error tolerance $\varepsilon>0$. The attainment of weak recovery is a crucial turning point; it allows the model to extract a meaningful signal of $\beta$ from $y$, significantly reducing the data requirements for the subsequent phase. Specifically, once $\left\langle\beta,{\bm{u}}^{(1)}\right\rangle>1/\mathrm{polylog}(d)$ is established, the signal strength becomes $O(1/\mathrm{polylog}(d))$, which is independent of $\mathrm{ge}(\sigma_{*}^{\mathrm{test}})$. This independence ensures that the sample size needed for strong recovery remains unaffected by $\mathrm{ge}(\sigma_{*}^{\mathrm{test}})$. Moreover, our approach achieves superior sample complexity compared to the prior work by Lee et al. (2024). While their work suggested a linear increase in $\left\langle\beta,{\bm{u}}^{(n)}\right\rangle$, which resulted in the required data length $N=\Theta_{\varepsilon}(\varepsilon^{-2})$, we find that beyond a certain point, the error $1-\left\langle\beta,{\bm{u}}^{(n)}\right\rangle$ converges to 0 geometrically. This acceleration yields an improved sample complexity of $N=\Theta_{\varepsilon}(\frac{1}{\varepsilon}\log{\frac{1}{\varepsilon}})$. See Appendix E for further details.